OpenStudy (anonymous):

Jessica is standing 20 feet away from a skyscraper in New York when she measures the angle of elevation to the top of the building as 88.8°. To the nearest foot, how tall is the building?

OpenStudy (anonymous):

sin88.8=height/20

OpenStudy (anonymous):

inverse sin is used to find answer

OpenStudy (anonymous):

But I did that. And these are the answer choices: 20 ft 477 ft 48 ft 955 ft

OpenStudy (anonymous):

I'd use tan rather than sin. The 20 ft refers to the base of a right-angled triangle, called the adjacent side. You want to find the 'opposite' side. The formula for that is \[tan\theta = \frac{opposite}{adjacent}\]

OpenStudy (anonymous):

tan of the angle, remember SOH CAH TOA, Sin is opposite Over Hypotenuse, Cosine is Adjacent over Hypotenuse, Tangent is Opposite over adjacent. So the height (Opposite side to the angle) over the distance she is from the tower(Adjacent side to the angle) will equal the tangent of the angle.

OpenStudy (anonymous):

I got 22.1.../:

OpenStudy (anonymous):

What did you get for \(tan\theta\) first of all?

OpenStudy (anonymous):

If you're using a calculator, make sure it's set to degrees mode.

OpenStudy (anonymous):

OHHH! There's my problem, it worked now. I got 954.7. And rounding to the nearest foot, 955ft.

OpenStudy (anonymous):

Congrats!